# Context-Driven Performance Modeling for Causal Inference Operators on Neural Processing Units

### v1

@@ Abstract
The proliferation of large language models (LLMs) has driven demand for long-context inference on resource-constrained edge devices. However, deploying these models on Neural Processing Units (NPUs) presents significant challenges due to the architectural mismatch: quadratic complexity of standard attention mechanisms conflicts with memory and compute patterns of edge accelerators. This paper presents a comprehensive performance analysis of various causal inference operators on a modern NPU. We benchmark standard quadratic attention against several sub-quadratic alternatives, including structured state-space and linear attention models. Our analysis reveals that while sub-quadratic methods offer superior scalability, they introduce distinct computational bottlenecks on the NPU’s specialized execution units. We identify that quadratic attention becomes severely memory-bound, suffering from cache inefficiency and pipeline stalls exceeding 95% at long contexts. In contrast, sub-quadratic models can become compute-bound on programmable vector cores. These findings provide critical insights for the co-design of hardware-aware models and optimization strategies to enable on-device AI inference with long-contexts.
@@ I Introduction
Long-context inference has become essential for document understanding, conversational AI, and real-time decision systems [1, 2]. Applications from medical record analysis to legal contract review increasingly demand processing of 100K+ token sequences. While cloud solutions can handle these workloads, privacy concerns, latency requirements, and operational costs are driving deployment to edge devices. Modern edge platforms now integrate Neural Processing Units (NPUs)—specialized accelerators that deliver exceptional efficiency for data-parallel operations but face severe constraints: limited scratchpad memory (typically 2-4MB), systolic execution patterns, and strict power budgets.
Transformer-based models like Llama deliver state-of-the-art quality but require quadratic computation and linear memory growth with context length [3]. At just 16K tokens, the Key-Value cache consumes over 768MB—more than 30x the capacity of leading NPUs. State-space models (SSMs) like Mamba offer linear scaling but underutilize NPU parallelism due to their recurrent nature during autoregressive decoding inference [4]. This fragmentation forces practitioners into hardware-specific compromises, limiting adoption of long-context capabilities where they’re needed most: on-device.
Recent works have looked into hybrid architectures or ways towards creating a general class of operators for causal inference [5, 6, 7]. While these efforts advance the theoretical understanding, a critical gap remains in characterizing how these diverse operators perform on real-world, resource-constrained hardware like NPUs. The architectural trade-offs made by these models—such as prioritizing data locality, computational regularity, or state compression—have profound, yet unquantified, implications for on-device performance. This paper bridges that gap by providing a comprehensive performance analysis and modeling of causal inference operators on a modern NPU. We move beyond theoretical complexity and provide a rigorous empirical study of latency, throughput, and hardware utilization, culminating in a Roofline analysis that explains the performance landscape.
@@ I-A Our Contributions
  • A comprehensive performance benchmark of quadratic and sub-quadratic attention operators on a real-world NPU, analyzing latency, throughput, and pipeline efficiency.
@@ II-A Fundamental Tradeoff for Long-Context Inference
We introduce this tradeoff not only to compare model classes, but also because it forms the architectural basis for the structured operators analyzed in this paper. Many recent causal inference variants—such as Toeplitz, Retentive, and Fourier—draw inspiration from state-space models, embedding recurrent or convolutional priors directly into attention layers [8, 4, 5, 9, 10]. These hybrid mechanisms blend the long-range capacity of traditional attention with the computational efficiency and inductive structure of SSMs, making the memory-state tradeoff a key element for evaluating modern causal operators on resource-constrained hardware platforms such as NPU.
@@ II-B Heterogeneous Edge Platform with Neural Processing Units (NPU)
State-of-the-art heterogeneous edge platforms—such as Intel Core Ultra processors (Intel AI PC)—combine traditional multi-core CPUs with specialized accelerators, including Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These diverse compute elements are integrated into a single System-on-Chip (SoC) architecture, featuring a unified system memory that facilitates seamless communication and efficient data sharing across all cores.
@@
Figure 2: NPU dataflow architecture with processing elements (PEs) and accumulator hierarchy. Note the absence of high-bandwidth memory for persistent state storage.
@@
Each of the three compute cores is designed to exploit specific types of workloads, leveraging their distinct system architectures and local memory hierarchies. The classic multi-core CPU, equipped with a three-level cache hierarchy, excels at handling general-purpose, logic-intensive workloads. GPUs feature Xe cores, each comprising multiple vector engines that are optimized for high data parallelism. Every Xe core includes a Shared Local Memory (SLM) block, accessible by all its vector engines to facilitate efficient intra-core communication. In NPUs, compute acceleration is driven by the Data Path Unit (DPU), comprising of a PE array which employs a structured systolic array architecture to perform operations such as matrix multiplication with minimal data movement. NPUs also incorporate Streaming Hybrid Architecture Vector Engines (SHAVE), which support parallel execution of general-purpose tasks and activation function engine to efficiently compute activations. To further enhance efficiency, Direct Memory Access (DMA) engines are integrated into the NPU, enabling high-throughput data transfers from the global shared system memory to the local, software-managed cache. Due to their controlled data movement and predictable, statically scheduled execution flow, NPUs achieve higher energy efficiency for AI workloads compared to GPUs, which involve complex, dynamic execution patterns and data transfers. This energy efficiency is a critical requirement in resource-constrained edge computing environments, where power limitation is a significant design consideration.
II-C Causal Operators for Attention
Figure 3: From left to right, top to bottom are Full Causal, Toeplitz, Fourier, Retentive Decay, Semiseparable, and Linear Structured attention masks.
@@ II-C Causal Operators for Attention
We describe in further detail which Causal Attention Masks we select to comprehensively survey and perform performance model analysis of a more general class of causal matrices seen in Figure 3.
@@
Linear Attention: Replaces the softmax with kernel functions to enable linear complexity:
$\text{LinearAttention}(Q,K,V)=\phi(Q)(\phi(K))^{T}V$
where $\phi$ is a kernel function (in our implementation, low-rank projections). This reduces complexity from $O(N^{2})$ to $O(N)$.
Toeplitz Structured Attention [11]: Constrains the attention matrix to have constant diagonals:
$W_{ij}=\gamma^{|i-j|}$
$\text{ToeplitzAttention}=\text{softmax}(QK^{T}\odot W)V$
where $\gamma$ is a decay factor. This models position-based decay patterns.
Fourier Structured Attention [12]: Leverages the convolution theorem to compute attention in frequency domain:
@@
Retentive Attention [8]: Introduces a decay mechanism that assigns exponentially decreasing weights to tokens based on relative position:
@@
where $\gamma\in(0,1)$ is the decay rate. This maintains causality while modeling recency bias in attention weights. Retentive attention provides tunable context compression and exhibits hardware-friendly diagonal structure.
@@ III-A Experimental Setup
We conduct a series of microbenchmarks to evaluate the performance of various attention mechanisms on a Neural Processing Unit (NPU). The objective is to analyze device utilization and latency as a function of context length. All experiments are executed on a system with the hardware specifications detailed in Table I. The NPU integrates a Digital Signal Processor (DSP) for control flow, a Data Path Unit (DPU) with a systolic array for matrix multiplication, programmable SIMD vector processors (SHAVE cores), and a high-bandwidth Direct Memory Access (DMA) engine.
@@
CPU 16 cores (8P + 8E) Control Logic
NPU 10 TOPS @ 35W Systolic Array Acceleration
DPU (PE Array) $\times128$ INT8 Matrix Multiplication
@@
DMA Bandwidth 64 GB/s Data Movement
@@
Memory 32 GB LPDDR5X Global Buffer
III-B Device Utilization Analysis

### v2

@@ Abstract
The proliferation of large language models has driven demand for long-context inference on resource-constrained edge platforms. However, deploying these models on Neural Processing Units (NPUs) presents significant challenges due to architectural mismatch: the quadratic complexity of standard attention conflicts with NPU memory and compute patterns. This paper presents a comprehensive performance analysis of causal inference operators on a modern NPU, benchmarking quadratic attention against sub-quadratic alternatives including structured state-space models and causal convolutions. Our analysis reveals a spectrum of critical bottlenecks: quadratic attention becomes severely memory-bound with catastrophic cache inefficiency, while sub-quadratic variants span from compute-bound on programmable vector cores to memory-bound by data movement. These findings provide essential insights for co-designing hardware-aware models and optimization strategies to enable efficient long-context inference on edge platforms.
@@ I Introduction
Long-context inference has become essential for document understanding, conversational AI, and real-time decision systems [chung2025long, qian2024long]. Applications from medical record analysis to legal contract review increasingly demand processing of 100K+ token sequences. While cloud solutions can handle these workloads, privacy concerns, latency requirements, and operational costs are driving deployment to edge devices. Modern edge platforms now integrate Neural Processing Units (NPUs)—specialized accelerators that deliver exceptional efficiency for data-parallel operations but face severe constraints: limited scratchpad memory (typically 2-4MB), spatial dataflow execution patterns, and strict power budgets.
Transformer-based models like Llama deliver state-of-the-art quality but require quadratic computation and linear memory growth with context length [dao2022flash]. At just 16K tokens, the Key-Value cache consumes over 768MB—more than 30x the capacity of leading NPUs. State-space models (SSMs) like Mamba offer linear scaling but underutilize NPU parallelism due to their recurrent nature during autoregressive decoding inference [gu2024mamba]. This fragmentation forces practitioners into hardware-specific compromises, limiting adoption of long-context capabilities where they’re needed most: on-device.
Recent works have looked into hybrid architectures or ways towards creating a general class of operators for causal inference [dao2024ssd, lieber2024jambahybridtransformermambalanguage, dong2024hymbahybridheadarchitecturesmall]. While these efforts advance the theoretical understanding, a critical gap remains in characterizing how these diverse operators perform on real-world, resource-constrained hardware like NPUs [hao_llm_perf, rak_part_2]. The architectural trade-offs made by these models—such as prioritizing data locality, computational regularity, or state compression—have profound, yet unquantified, implications for on-device performance. This paper bridges that gap by providing a comprehensive performance analysis and modeling of causal inference operators on a modern NPU. We move beyond theoretical complexity and provide a rigorous empirical study of latency, throughput, and hardware utilization, culminating in a Roofline analysis that explains the performance landscape.
@@ I-A Our Contributions
  • A comprehensive performance benchmark of quadratic and sub-quadratic attention operators on a real-world edge NPU, analyzing latency, throughput, and pipeline efficiency.
@@ II-A Fundamental Tradeoff for Long-Context Inference
We introduce this tradeoff not only to compare model classes, but also because it forms the architectural basis for the structured operators analyzed in this paper. Many recent causal inference variants—such as Toeplitz, Retentive, and Fourier—draw inspiration from state-space models, embedding recurrent or convolutional priors directly into attention layers [sun2023retentivenetworksuccessortransformer, gu2024mamba, dao2024ssd, poli2023hyena, peng2023rwkv]. These hybrid mechanisms blend the long-range capacity of traditional attention with the computational efficiency and inductive structure of SSMs, making the memory-state tradeoff a key element for evaluating modern causal operators on resource-constrained compute units such as NPU.
@@ II-B Heterogeneous Edge Platform with Neural Processing Units (NPU)
State-of-the-art heterogeneous edge platforms—such as Intel® Core™ Ultra processors (Intel AI PC)—combine traditional multi-core CPUs with specialized accelerators, including Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These diverse compute units are integrated into a single System-on-Chip (SoC) architecture, featuring a unified system memory that facilitates seamless communication and efficient data sharing across all cores. Representative commercial edge-AI platforms include AMD’s Ryzen AI Processors [AMD_Ryzen_AI_2025], Qualcomm’s Hexagon NPU [QualcommHexagon2025], and Google’s Coral Edge TPU/NPU family [Rutledge2025CoralNPU], which collectively exemplify the growing diversity of on-device neural accelerators across CPU–GPU–NPU SoCs. Recent benchmarking studies [rak_bench_aiml, pal_bench_aiml, zuo_dhru] further characterize these heterogeneous systems in terms of performance–energy tradeoffs for ML workloads.
@@
Figure 2: NPU dataflow architecture with processing elements (PEs) and accumulator hierarchy. Note the absence of high-bandwidth memory for persistent context storage.
@@
Each of the three compute units is designed to exploit specific types of workloads, leveraging their distinct system architectures and local memory hierarchies. The classic multi-core CPU, equipped with a three-level cache hierarchy, excels at handling general-purpose, logic-intensive workloads. GPUs feature Xe cores, each comprising multiple vector engines that are optimized for high data parallelism. Every Xe core includes a Shared Local Memory (SLM) block, accessible by all its vector engines to facilitate efficient intra-core communication. In NPUs, compute acceleration is driven by the Data Path Unit (DPU), comprising of a PE array which employs a structured spatial MAC array architecture to perform operations such as matrix multiplication with minimal data movement. NPUs also incorporate Streaming Hybrid Architecture Vector Engines (SHAVE), which support parallel execution of general-purpose tasks and activation function engine to efficiently compute activations. To further enhance efficiency, Direct Memory Access (DMA) engines are integrated into the NPU, enabling high-throughput data transfers from the global shared system memory to the local, software-managed cache. Due to their controlled data movement and predictable, statically scheduled execution flow, NPUs achieve higher energy efficiency for AI workloads compared to GPUs, which involve complex, dynamic execution patterns and data transfers. This energy efficiency is a critical requirement in resource-constrained edge computing environments, where power limitation is a significant design consideration.
II-C Causal Operators
Across operator types, causality is preserved in different ways. In Figure 3, we show how each type of operator maintains causality—the fundamental requirement that position $i$ cannot access information from future positions $j>i$.
Figure 3: Preserving causality across operator/context types.
Attention-based Causality: Causal attention mechanisms enforce temporal ordering through explicit masking of the attention matrix. The lower-triangular structure ensures that each query position can only attend to key-value pairs from current and previous positions. This is achieved by setting future attention weights to -$\infty$ before the softmax operation, effectively zeroing out their contribution. While this provides maximum expressiveness—each position has access to all preceding context—it results in a dense $N\times N$ attention matrix with $O(N^{2})$ computational complexity.
SSM-based Causality: State-space models maintain causality through sequential state updates. At each timestep $t$, the model receives input $x_{t}$, updates the hidden state $h_{t}$ based only on the previous state $h_{t-1}$ and current input, then generates output $y_{t}$. This recurrent formulation inherently prevents information leakage from future timesteps, as the state evolution is strictly unidirectional. The fixed-size hidden state $h_{t}\in\mathbb{R}^{d_{s}}$ compresses all past context into $O(d_{s})$ memory, trading expressiveness for memory efficiency. During inference, this sequential dependency limits parallelization but enables constant memory footprint.
Convolution-based Causality: Causal convolutions preserve temporal ordering through asymmetric padding and kernel design. The convolution kernel $w\in\mathbb{R}^{K}$ only spans past positions: output at position $t$ is computed as a weighted sum over positions $\{t,t-1,...,t-K+1\}$. This is implemented via left-side padding that adds K1K-1 zeros before the sequence, ensuring that the receptive field never extends beyond the current position. Unlike attention’s global context or SSMs’ compressed state, convolutions provide a sliding window view with local receptive fields, offering $O(NK)$ complexity with tunable context length $K$.
The key architectural distinction lies in how context is accessed: attention maintains explicit key-value memories, SSMs compress context into fixed states, and convolutions aggregate through local kernels. These design choices create fundamentally different memory-compute tradeoffs that manifest distinctly on NPU hardware, as we analyze in subsequent sections.
II-C1 Attention-based Structured Masks
Figure 4: Structured masked attention variants with differing causal matrices.
@@ II-C1 Attention-based Structured Masks
We describe in detail the causal attention masks selected to comprehensively survey and perform performance model analysis of a general class of causal matrices. In Figure 4, we show sample structured masks which preserve causality in attention.
@@
Banded (Toeplitz) Structured Attention [qin2023toeplitzneuralnetworksequence]: Constrains the attention matrix to have constant diagonals with a fixed bandwidth $w$:
$W_{ij}=\begin{cases}\gamma^{|i-j|}&|i-j|\leq w\\ 0&\text{otherwise}\end{cases}$
$\text{BandedAttention}=\text{softmax}(QK^{T}\odot W)V$
where $\gamma$ is a decay factor and $w$ is the band width. This models position-based decay patterns while limiting the context window, reducing complexity to $O(Nw)$.
Semiseparable Attention [dao2024ssd]: Decomposes the attention matrix into low-rank plus diagonal structure, enabling efficient parallel algorithms:
$A=\text{tril}(PQ^{T})+D$
where $P,Q\in\mathbb{R}^{N\times r}$ are low-rank factors and $D$ is diagonal. This admits $O(Nr^{2})$ complexity through parallel prefix scans while maintaining full causal context.
Fourier Structured Attention [leethorp2022fnetmixingtokensfourier]: Leverages the convolution theorem to compute attention in frequency domain:
@@
Retentive Decay Attention [sun2023retentivenetworksuccessortransformer]: Introduces an exponential decay mechanism that assigns decreasing weights based on relative position:
@@
where $\gamma\in(0,1)$ is the decay rate. This maintains causality while modeling recency bias and exhibits hardware-friendly diagonal structure.
II-C2 State-Space Model-based Recurrence
State-space models with fixed-size persistent state memory take on different computational forms for training versus inference [gu2024mamba].
Sequential Recurrence Mode: Computes the state sequence sequentially, ideal for autoregressive decoding:
$h_{t}=Ah_{t-1}+Bx_{t}$
$y_{t}=Ch_{t}$
where $A\in\mathbb{R}^{d_{s}\times d_{s}}$, $B\in\mathbb{R}^{d_{s}\times d_{m}}$, $C\in\mathbb{R}^{d_{m}\times d_{s}}$. This has $O(n)$ time complexity with $O(d_{s})$ memory but inherent sequential dependencies.
Parallel (Scan) Mode: Computes the state sequence in parallel using associative prefix sums [dao2024ssd]:
$h_{t}=\sum_{k=0}^{t}A^{t-k}Bx_{k}$
$y=\text{parallel\_scan}(\{(A,Bx_{i})\}_{i=1}^{N})\times C$
This leverages parallel prefix algorithms for $O(\log n)$ depth, enabling efficient training.
Chunked Recurrence Mode: Hybrid approach that processes input in chunks of size $L$, computing intra-chunk representations in parallel while maintaining recurrent state across chunks:
$h_{t+L}=A^{L}h_{t}+\sum_{i=0}^{L-1}A^{L-1-i}Bx_{t+i}$
This balances parallelism within chunks ($O(L)$ parallel operations) with sequential cross-chunk dependencies, providing a favorable memory-compute tradeoff for long sequences.
II-C3 1D Convolution-based Models
Convolutional approaches to sequence modeling apply causal filters that respect temporal ordering [poli2023hyena].
Direct Convolution: Explicitly computes the causal convolution:
$y_{t}=\sum_{k=0}^{K-1}w_{k}x_{t-k}$
where $w\in\mathbb{R}^{K}$ is the learned kernel. Complexity is $O(NK)$ for sequence length $N$ and kernel size $K$. Hardware-friendly due to regular memory access patterns.
FFT Convolution: Exploits the convolution theorem to compute in frequency domain:
$y=\mathcal{F}^{-1}(\mathcal{F}(x)\odot\mathcal{F}(w))$
Achieves $O(N\log N)$ complexity but requires special handling to maintain causality through appropriate zero-padding. Efficiency depends on FFT library optimizations.
Dilated Convolution: Expands receptive field through exponentially-spaced sampling:
$y_{t}=\sum_{k=0}^{K-1}w_{k}x_{t-d^{k}}$
where $d$ is the dilation rate. Achieves receptive field of $O(d^{K})$ with only $K$ parameters, enabling efficient long-range modeling with $O(NK)$ operations.
@@ III-A Experimental Setup
We conduct a series of microbenchmarks to evaluate the performance of various attention mechanisms on a Neural Processing Unit (NPU). The objective is to analyze device utilization and latency as a function of context length. All experiments are executed on a system with the hardware specifications detailed in Table I. The NPU integrates a Digital Signal Processor (DSP) for control flow, a Data Path Unit (DPU) with a spatial MAC array for matrix multiplication, programmable SIMD vector processors (SHAVE cores), and a high-bandwidth Direct Memory Access (DMA) engine.
III-B Performance Instrumentation Methodology
To characterize the fine-grained execution behavior of causal operators on NPU hardware, we employ a hierarchical instrumentation strategy that captures both aggregate performance metrics and detailed microarchitectural events. Our methodology bridges the gap between high-level operator semantics and low-level hardware resource utilization, enabling rigorous bottleneck identification.
At the coarse-grained level, we collect per-layer performance counters that attribute execution time to specific hardware units. Each neural network layer is instrumented to record its execution duration, the hardware unit responsible for computation (DPU spatial MAC array, SHAVE vector cores, or DMA engine), and its execution status. This attribution is critical because NPU performance is fundamentally heterogeneous: matrix multiplications execute on the DPU with predictable spatial dataflow, while element-wise operations such as softmax and exponential decay masks are dispatched to SHAVE cores with irregular memory access patterns. By tracking which hardware unit executes each operation, we can identify architectural mismatches where operators stress units ill-suited to their computational structure.
To understand temporal behavior beyond aggregate metrics, we instrument the NPU runtime with fine-grained event tracing. This captures pipeline stage boundaries—specifically the push (data movement to NPU), pull (result retrieval), and initialization phases—allowing us to decompose total latency into constituent components. Memory allocation events are correlated with performance degradation, revealing when the 4 MB scratchpad overflows and triggers costly evictions to system memory. Compilation metadata, including estimated inference latency from the static analyzer and peak memory consumption during graph optimization, provides ground truth for comparing predicted versus observed performance. These traces are timestamped at microsecond resolution, enabling precise reconstruction of execution timelines and identification of pipeline stalls.
Our experimental protocol isolates steady-state performance from transient compilation effects. For each operator-context pair, we execute three warmup inferences to populate instruction caches and finalize just-in-time compilation, then measure performance over ten timed runs for CPU and GPU (reduced to one for NPU due to its longer absolute latency). This repetition accounts for variance in system noise while maintaining practical evaluation time. Performance counters are collected on the final inference to avoid measurement perturbation, while execution traces are captured in dedicated runs with enhanced logging to preserve timing fidelity. This two-phase approach—coarse metrics for statistical characterization and fine-grained traces for causal analysis—provides the empirical foundation for our roofline model and bottleneck attribution.
III-C NPU Compiler Optimizations and Tensor Layout
The NPU compiler plays a critical role in bridging the semantic gap between high-level operator graphs and low-level hardware primitives. A fundamental transformation performed during compilation is the enforcement of canonical tensor layouts that match the NPU’s dataflow architecture. Specifically, the compiler ensures all tensor operands conform to the NCHW (batch, channel, height, width) memory layout, which aligns with the spatial batched 2D convolution primitive that underlies all DPU matrix operations.
This layout standardization is essential because the DPU’s spatial MAC array architecture is optimized for regular, strided memory access patterns characteristic of spatial convolutions. By representing even non-convolutional operations—such as matrix multiplication in attention mechanisms—as spatial convolutions with unit kernel size, the compiler enables these operations to leverage the DPU’s 128×128 processing element array with minimal data movement. The NCHW format facilitates this mapping: the height and width dimensions are treated as spatial coordinates for spatial dataflow execution, while channels are distributed across the accumulator hierarchy to exploit data reuse.
However, this compilation strategy introduces constraints that impact certain operator classes. Operators requiring complex tensor reshaping or transposition—such as the frequency-domain computations in Fourier attention or the diagonal indexing in Toeplitz masks—incur additional overhead as the compiler must insert explicit data layout transformations. These transformations manifest as DMA operations that shuffle data between incompatible formats, contributing to the DMA-bound behavior observed in our profiling results. Furthermore, operations that do not naturally map to the 2D spatial convolution primitive, such as the element-wise multiplication in retentive decay masks, are offloaded to SHAVE cores where the lack of spatial dataflow reduces parallelism.
The compiler also performs static analysis to estimate resource requirements and schedule operations across the NPU’s heterogeneous units. This includes computing an activity factor—the percentage of processing elements actively computing versus idle—and estimating inference latency based on the critical path through the dependency graph. These estimates, which we extract from compilation logs, provide a theoretical performance upper bound that we compare against measured throughput to quantify the gap between ideal and achieved hardware utilization. The discrepancy, often exceeding 20× for poorly-suited operators, reveals fundamental architectural mismatches that cannot be resolved through software optimization alone and motivate the co-design insights presented in our discussion.
@@ III-C NPU Compiler Optimizations and Tensor Layout
CPU Intel® Core™ Ultra 9 185H Control Logic
16 cores (8P + 8E)
NPU 10 TOPS @ 35W Spatial MAC Array Acceleration
DPU (PE Array) 128×128 INT8 Matrix Multiplication
@@
Bandwidth 182 GB/s Data Movement
@@
Memory 64 GB LPDDR5X Global Buffer
III-D Operator Decomposition and Hardware Unit Mapping
A critical insight for understanding NPU performance is that high-level causal operators decompose into distinct primitive operations, each with deterministic hardware unit assignments. This decomposition creates predictable bottleneck patterns: operators become memory-bound, compute-bound, or vector-bound depending on their constituent primitive mix. By analyzing the operator graph prior to execution through static analysis of the OpenVINO computation graph, we can forecast which hardware unit will limit throughput—a capability that distinguishes NPUs from dynamically-scheduled GPUs.
Table II categorizes primitive operations by their target hardware unit. Data movement operations—transpose, reshape, concatenation, slicing—are serviced by the DMA engine, which orchestrates transfers between system memory and the 4 MB on-chip scratchpad at up to 182 GB/s bandwidth. Matrix multiplication operations, including GEMM and batched convolution, execute on the DPU’s 128×128 spatial MAC array, achieving peak efficiency on dense, regularly-strided NCHW tensors. Element-wise operations—activation functions, scaling, masking, and FFT computations—are offloaded to the eight 1.4 GHz SHAVE vector cores, which lack the data reuse locality of spatial dataflow execution.
TABLE II: Hardware unit mapping for primitive operations. Assignment is deterministic and performed during static compilation, enabling predictive bottleneck analysis.
Hardware Unit Operation Type Examples
DMA Data Movement Transpose, Reshape, Concat, Slice
DPU Matrix Math GEMM, GEMV, Batched Conv
SHAVE Element-wise Softmax, Multiply, Divide, FFT
Through systematic decomposition of operator computation graphs, we quantify the exact primitive operation counts for each causal operator variant. Figure 5 presents this analysis for three representative operators exhibiting distinct architectural profiles. The decomposition reveals stark imbalances that directly predict measured performance bottlenecks.
Figure 5: Primitive operation distribution across NPU hardware units for three representative causal operators.
Conv1D FFT represents the pathological case of DMA saturation: 10 of its 16 operations (62%) are data movement primitives—four transpose operations to convert between NCHW and NCL layouts, two complex tensor constructions via concatenation, and additional reshape/slice operations for FFT compatibility. With only 2 DPU operations (12%) for dimensional projections, the spatial MAC array remains starved while the DMA engine saturates transferring 64 GB/s of layout transformations. This explains the measured 0.34 GOP/s performance despite 15 Ops/Byte arithmetic intensity, confirming that memory access patterns—not FLOP counts—dominate NPU execution.
Toeplitz Banded Attention achieves the most DPU-favorable distribution with 7 of 10 operations (70%) mapping to matrix multiplication: three QKV projections, KTV computation, 2D convolution (the core Toeplitz operation), Q-convolution product, and output projection. The three DMA operations (30%) handle unavoidable tensor reshaping for convolution setup. Critically, Toeplitz contains zero SHAVE operations—decay patterns are encoded directly into the convolution kernel structure, eliminating the vector core bottleneck entirely. This architectural alignment results in 87.9% cache efficiency and only 36.4% pipeline stalls at $N=4096$ (Table X), achieving 3.5× higher utilization than quadratic attention despite similar arithmetic intensity.
Fourier Structured Attention exhibits a deceptively balanced distribution: 5 DMA operations (36%), 4 DPU operations (29%), and 5 SHAVE operations (36%). However, this balance creates architectural ping-ponging—the execution graph alternates between DMA complex tensor construction, SHAVE DFT/IDFT transforms, and DPU linear projections, forcing frequent cross-unit synchronization. Each DFT requires three DMA operations (to/from complex format plus result extraction) that bracket SHAVE computation, creating pipeline bubbles where the spatial MAC array idles awaiting frequency-domain results. This explains Fourier’s bifurcated behavior: DPU-bound at short contexts when DFT overhead is amortized, transitioning to DMA-bound beyond $N>512$ tokens when tensor management saturates the 64 GB/s bandwidth.
The predictability of these bottlenecks arises from the NPU’s static compilation model. During graph optimization, the compiler deterministically assigns each node to a hardware unit: operations requiring tensor layout transformations route to DMA (transpose, reshape, slice), dense linear algebra routes to DPU (matmul, conv2d), and non-fused element-wise operations route to SHAVE (softmax, multiply, FFT). By counting the number and computational cost of each primitive type in an operator’s graph, we estimate execution time distribution across units—predictions that match measured utilization within 5% for all operators tested.
This decomposition analysis reveals a fundamental design principle for NPU-efficient operators: minimize cross-unit dependencies while maximizing DPU utilization. Toeplitz’s success stems from fusing multiple logical operations into single DPU-executable convolutions that avoid SHAVE offloading. Fourier’s mediocrity results from unavoidable cross-unit coordination between frequency-domain computation and spatial projections. Conv1D FFT’s failure exemplifies the catastrophic impact of DMA saturation when data layout transformations dominate over arithmetic. These insights—derived from static graph analysis and validated by runtime profiling—provide actionable guidance for co-designing hardware-aware causal operators.
III-E Device Utilization Analysis